OpenStudy (anonymous):

check answers! k(0)=x+7 k(0)=-7+7 k(-10)=x+7=

OpenStudy (anonymous):

k(0)=x+7 = 7 k(0)=-7+7 = 0 k(-10)=x+7= -3